A bag contains 7 red marbles, 6 blue marbles, and 8 green marbles. A randomly drawn marble is blue. O on la
Answer #1

SOLUTION :


Total marbles in the bag = 7 + 6 + 8 = 21 


P( blue) 

= blue marbles in the bag  / Total marbles in the bag

= 6/21 = 2/7 = 3rd option  (ANSWER)

answered by: Tulsiram Garg